OpenStudy (anonymous):

Please help!!!! Which tangent line gives a better approximation x=1? f(x)=-x^(3)-2x^(2)+x+1 and g(x)=ln(x+1)+1?

OpenStudy (anonymous):

The equation for the first one is 3x^(2)-4x+1 and the second is (1/x+1).

OpenStudy (anonymous):

The tangent line equation for the first one is y=x+1 and it is the same for the second one y=x+1.

OpenStudy (anonymous):

How do I see which tangent line gives a better approximation at x=1? Thanks!

OpenStudy (amistre64):

f(x)=-x^(3)-2x^(2)+x+1 and g(x)=ln(x+1)+1 f'(x)=-3x^(2)-4x+1 and g'(x)=1/(x+1) f'(1) = -6 g'(1) = 1/2 im not sure how you can to the conclusion that they have the same tangent lines since the clearly have different slopes

OpenStudy (anonymous):

The question part b said "Show that g(x) has the same tangent line as f(x) at x=0".

OpenStudy (amistre64):

ah, at x=0 ...

OpenStudy (anonymous):

Yeah, both at x=0.

OpenStudy (amistre64):

f(0)=1 and g(0)=1 and the derivatives have the same slope at x=0 tanf = tang = x+1

OpenStudy (amistre64):

at x=1 we have a linear approximation of 2 right?

OpenStudy (anonymous):

Yeah!

OpenStudy (amistre64):

now we just compare this to the actual values to see which is the better approximation

OpenStudy (amistre64):

what is f(1) and g(1)

OpenStudy (anonymous):

f(1)= -1

OpenStudy (amistre64):

so at some error: tanf(1) + error = f(1) 2 + error = -1 error = -3, of as a distance, we are off by 3 compare tang(1) + error = g(1) and define the error for it

OpenStudy (anonymous):

Does that mean that it's not a good approx.?

OpenStudy (amistre64):

dunno, we are seeing which function it approximates better, so the one with the smallest error wins

OpenStudy (anonymous):

I thought somewhere we had to use f(a)+f'(a)(x-a) for linear approximations?

OpenStudy (amistre64):

we did

OpenStudy (amistre64):

thats how we made the tangent lines

OpenStudy (amistre64):

tanf = f(0)+f'(0)(x-0) tang = g(0)+g'(0)(x-0)

OpenStudy (anonymous):

Oh okay. So at x=1 it is supposed to be close to 2?

OpenStudy (amistre64):

no, at x=1, the approximation of the tangent line is 2 the functions will be whatever they are going to be. linear approximations are only so good for a limited interval ... at least as far as an error rate is concerned

OpenStudy (amistre64):

all we are doing in this is seeing how which function has the better approximation is all

OpenStudy (anonymous):

I'm sorry. I was really confused!

OpenStudy (amistre64):

does it makes sense yet?

OpenStudy (amistre64):

at x=0, everything is great .... as we move away from 0, the curves bend away from the line. one curves further away than the other.

OpenStudy (anonymous):

could I graph this on a calculator to see?

OpenStudy (anonymous):

So the ln(x+1)+1 function better approximates at x=1?

OpenStudy (amistre64):

the tangent line better approximates ln(x+1)+1 at x=1, yes it has a smaller error

OpenStudy (anonymous):

Okay! I get it now. Thank you. I think I will just graph it from now on to understand!

OpenStudy (amistre64):

if that helps :) good luck